# 22affiliationtext: Booth School of Business

### v1

@@ 4.2 Details
Our goal is to augment the decoder of a generative model (without modifying the internals of the existing decoder) in such a way to identify $d_{c}$ latent concepts from a black-box decoder. Moreover, we seek to accomplish this without sacrificing expressivity: The input layer to the decoder as well as the output layer of the encoder should both be arbitrary. This not only allows for arbitrary flexibility in both the latent dimension as well as the structure of the black-box encoder-decoder and its latent space, but also makes implementation significantly simpler: The learning of latent concepts is completely abstracted away from the encoding step.
To accomplish this, we assume given a black-box encoder-decoder pair, denoted by $\operatorname{enc}(\mathbf{x})$ and $\operatorname{dec}(\mathbf{z})$, respectively. We modify the decoder by appending a context module to the head of the decoder. This context module is divided into three layers: A representation layer, an intervention layer, and an expressive layer.
@@
    The first representation layer, as the name suggests, learns to represent concepts by implicitly inverting their linear representations $\mathbf{c}_{j}=C_{j}\mathbf{e}$ from the embeddings of the encoder. This is a linear layer between $\mathbf{c}\to\mathbf{e}$.
@@
    The intervention layer embeds these concept representations into an overall reduced form SEM as suggested by (1). Each concept corresponds to its own context where it has been intervened upon. A key part of this layer is how it can be used to learn and enforce interventional semantics through this shared SEM. This layer corresponds to the second layer $\boldsymbol{\varepsilon}\to\mathbf{c}$ in (2).
@@
Crucially, no part of this SEM is fixed or known—everything is trained end-to-end. In particular, we do not assume a known causal DAG or even a known causal order. This stands in contrast to previous work on causal generative models such as CausalGAN (Kocaoglu et al., 2017) and CausalVAE (Yang et al., 2021).
@@
For simplicity, assume here that the concepts are each one-dimensional with $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$; generalization to multi-dimensional concepts is more or less straightforward and explained below. The structural coefficients $\alpha_{kj}\in\mathbb{R}$ capture direct causal effects between concepts, with $\alpha_{kj}\neq 0$ indicating the presence of an edge $\mathbf{c}_{k}\to\mathbf{c}_{j}$. An intervention on the $j$th concept requires deleting all these incoming edges; i.e., setting $\alpha_{\cdot j}=0$, updating the outgoing edges from $\boldsymbol{\varepsilon}_{j}$, as well as replacing $\boldsymbol{\varepsilon}_{j}$ with a new $\boldsymbol{\varepsilon}_{j}^{\prime}$, which results from updating the incoming edges to $\boldsymbol{\varepsilon}_{j}$ in the preceding layer. Thus, during training, when we intervene on $\mathbf{c}_{j}$, we zero out the row $\alpha_{\cdot j}$ while replacing the column $\alpha_{j\cdot}$ with a new column $\beta_{j\cdot}$ that is trained specifically for this interventional setting. The result is $d_{c}+1$ intervention-specific layers: $A_{0}$ for the observational setting, and $A_{1},\ldots,A_{d_{c}}$ for each interventional setting. This is captured as a three-way tensor where each slice of this tensor captures a different interventional context that corresponds to intervening on different concepts. At inference time, to generate interventional samples, we simply swap out the $A_{0}$ for $A_{j}$. Moreover, multi-target interventions can be sampled by zero-ing out multiple rows and substituting multiple $\beta_{j\cdot}$’s and $\boldsymbol{\varepsilon}_{j}$’s.
@@
In the causal model described above, we assumed $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$ for simplicity. Everything carries through if we allow additional flexibility with $\dim(\mathbf{c}_{j})>1$ and $\dim(\boldsymbol{\varepsilon}_{j})>1$, potentially even with $\dim(\mathbf{c}_{j})\neq\dim(\boldsymbol{\varepsilon}_{j})$. In practice, we implement this via two width parameters $0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ and $0pt_{c}=\dim(\mathbf{c}_{j})$. The design choice of enforcing uniform dimensionality is not necessary, but is made here since in our experiments there did not seem to be substantial advantages to choosing nonuniform widths.
With these modifications, $\alpha_{kj}$ becomes a $0pt_{\varepsilon}\times 0pt_{c}$ matrix; instead of substituting rows and columns above, we now substitute slices in the obvious way. The three-way tensor $A$ is now $(d_{c}+1)\times 0pt_{\varepsilon}d_{c}\times 0pt_{c}d_{c}$-dimensional.
@@
An appealing aspect of this architecture is that under certain assumptions, it can be viewed as an approximation to an identifiable model over disentangled concepts. Formally, we assume given observations $\mathbf{x}$, concepts $\mathbf{c}$, and a collection of embeddings $\mathbf{e}$ such that $\mathbf{x}=f(\mathbf{e})$. Additionally assume a noisy version of the linear representation hypothesis, i.e., $\mathbf{c}_{j}=C_{j}\mathbf{e}+\varepsilon_{j}$ with $\varepsilon_{j}\sim\mathcal{N}(0,\Omega_{j})$. Then we have the following identifiability result:
Theorem 4.1 (Identifiability of linear representations).
Assume that the rows of each $C_{j}$ are chosen from a linearly independent set and $f$ is injective and differentiable. Then, given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$.
We do not claim to identify a causal graph, only the latent distribution $p(\mathbf{c})$ over the concepts and its functional structure $C_{j}$ within the data-generating process. As opposed to solving a causal representation learning problem, the causal semantics used in this paper are simply being ported into the generative modeling framework to give solutions to the problem of learning latent distribution structure in a completely unsupervised fashion. Theorem 4.1 shows that the set problems where there exists a structured latent distribution to be identified is nonempty. Moreover, the structure of these identifiable distributions exactly fits our model architecture, meaning that set of problems in which the proposed method will approximate a fully disentangled latent space is nonempty. In other words, there exist data-modeling problems that are solvable by the novel approach introduced in this paper. The technical difficulty in the proof of Theorem 4.1 is analyzing the behaviour of interventions in this model; since our main focus is on methodological aspects and the technical details are fairly involved, these details are deferred to the appendix.
@@ 4.3 Summary of architecture
    There is no coupling between the embedding dimension and the number of concepts, or the complexity of the SEM. In particular, this architecture allows for arbitrary black-box encoder-decoder architectures to be used to learn embeddings, from which a causal model is then trained on top of. This can be trained end-to-end, fine-tuned after pre-training, or even frozen, allowing for substantial computational savings. In our experiments, we illustrate this with both standard convolutional networks as well as NVAE (Vahdat and Kautz, 2020), which is a deep hierarchical VAE with millions of parameters.
@@
    The intervention layer is a genuine causal model that provides rigourous causal semantics that allow for sampling from arbitrary concept interventions, including interventions that have not been seen during training. But note that we do not claim to learn the structural coefficients, only the reduced model which is sufficient for concept interventions and sampling.
@@
    The architecture is based on an approximation to an identifiable concept model (Theorem 4.1), which provides formal justification for the intervention layer as well as reproducibility assurances.
@@
    The context module is itself arbitrarily flexible, meaning that there is no risk of information loss in representing concepts with the embeddings $\mathbf{e}$. Of course, information loss is possible if we compress this layer too much (e.g., by choosing $d_{c}$, $k$, $0pt_{\varepsilon}$, or $0pt_{c}$ too small), but this is a design choice and not a constraint of the architecture itself.
As a consequence, the only tradeoff between representational capacity and causal semantics is design-based: The architecture itself imposes no such constraints. The causal model can be arbitrarily flexible and chosen independent of the black-box encoder-decoder pair, which is also allowed to be arbitrarily flexible.
5 Empirical results
We divide the presentation here into two main parts: small models on (semi-)synthetic datasets (Section 5.1) versus larger models on more challenging datasets (Section 5.2). We emphasize key findings here but include complete results and experimental details in the appendix.
5.1 Small models on (semi-)synthetic data
To systematically investigate the interplay between architectural design, learned representations, and OOD performance, we begin with small-scale experiments with controlled environments. This setting enables rapid iteration, comprehensive ablation studies, and access to ground-truth metrics unavailable in complex real-world data. We evaluate models on two benchmarks:
  • MNIST: We use Morpho-MNIST (Castro et al., 2019) to allow for different interventional contexts, such as making digits thicker or thinner.
  • quad: A novel semi-synthetic visual environment with disentangled latents (colour, shape, size, orientation) and explicit OOD compositionality challenges (e.g., generating unseen combinations of background quadrant colours and central object attributes).
MNIST
On MNIST, we see (Table 1) results of our module across different ablations. The first column (Context module) is our module incorporated into a lightweight convolutional VAE. Ablation 1 is an analogously lightweight standard implementation of $\beta$-VAE trained only on the observational context. Ablation 2 is $\beta$-VAE trained on the full pooled dataset. Ablation 3 is similar to the context module, except that it ignores the contexts and pools everything into a single context (like Ablation 2). These results indicate that the cost of the structured representation learned by the context module is only a slight degradation in the numerical metrics.
Table 1: Ablation on MNIST (200 epochs)
Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
Validation Metrics
Validation ELBO (bpd) ↓ 0.2945 (0.0091) 0.1832 (0.0007) 0.2362 (0.0053) 0.2886 (0.0103)
Reconstruction Loss (bpd) ↓ 0.2649 (0.0093) 0.1469 (0.0008) 0.1998 (0.0050) 0.2575 (0.0111)
Reconstructed Distribution Evaluation
MMD (obs) ↓ 0.0092 (0.0027) 0.0000 (0.0000) 0.0028 (0.0022) 0.0124 (0.0032)
MMD (ivn) ↓ 0.0939 (0.0331) 0.0303 (0.0107) 0.0430 (0.0145) 0.0536 (0.0154)
Generated Distribution Evaluation
MMD (obs) ↓ 0.0099 (0.0011) 0.0017 (0.0007) 0.0097 (0.0030) 0.0098 (0.0019)
MMD (ivn) ↓ 0.0298 (0.0069) 0.0142 (0.0038) 0.0183 (0.0034) 0.0188 (0.0034)
Figure 2: Example images obtained from the quad dataset. (top) Each subfigure corresponds to a different context, indicated by the title (obs = observational; the others indicate a single-node intervention). (bottom) Each subfigure corresponds to a different double-concept intervention context, indicated by the title. Double-concept interventions were not included in the training data, allowing for genuinely OOD evaluation.
Controlled simulations on quad
While evaluations on unsimulated data are the gold standard for evaluating models, simulated environments are valuable for conducting controlled experiments under given conditions. To provide a controllable, synthetic test bed for evaluating composition in a visual environment, we developed a simple semi-synthetic benchmark, visualized in Figure 2.
quad is a visual environment defined by 8 concepts:
@@ 5.1 Small models on (semi-)synthetic data
    quad1: The colour of the first quadrant;
@@
    quad2: The colour of the second quadrant;
@@
    quad3: The colour of the third quadrant;
@@
    quad4: The colour of the fourth quadrant;
  5. 5.
    size: The size of the center object;
  6. 6.
    orientation: The orientation (angle) of the center object;
  7. 7.
    object: The colour of the center object;
  8. 8.
With the exception of shape, which is discrete, the remaining concepts take values in $[0,1]$.
Table 2: Ablation on Quad
Metric / Evaluation Context module Ablation 2
Generated Sample Evaluation
MMD (obs) ↓ 0.0099 (0.0058) 0.0187 (0.0063)
MMD (ivn) ↓ 0.0149 (0.0032) 0.0275 (0.0038)
MMD (ood) ↓ 0.0273 (0.0021) 0.0297 (0.0019)
Figure 3 shows OOD samples from a lightweight VAE trained on this dataset. Table 2 shows numerical results comparing our module (incorporated into a lightweight convolutional VAE) to Ablation 2 from above (analogously lightweight $\beta$-VAE trained on the full pooled dataset) for sample generation across the observational and interventional data (same distribution seen during training) along with OOD data. The quad data allows evaluating the sampled OOD distribution (which our module composes from seen contexts) with a ground truth OOD distribution (unseen during training).
Figure 3: Example generated OOD images from a simple, lightweight VAE on the quad benchmark. This shows a proof-of-concept that the concept module successfully composes concepts using even simple architectures. For performant architectures, see the NVAE results in Figure 5.2.
OOD generation vs. reconstruction
There is a crucial difference between OOD reconstruction and OOD generation. By OOD generation, we mean the ability to conditionally generate and combine novel interventions. For example, suppose during training the model only sees small, red objects along with big, blue objects. At test time, we wish to generate OOD samples of big, red objects or small, blue objects.
The crucial difference is that a model may be capable of OOD reconstruction, but not OOD generation. For example, we can always attempt to reconstruct an OOD sample and evaluate its reconstruction error. This is a common metric that has been used in previous work to evaluate the OOD capabilities of generative models (e.g. Xu et al., 2022; Mo et al., 2024; Montero et al., 2022; Schott et al., 2022). But unless the model learns specific latent concepts corresponding to size or colour, then we have no control over the size or colour of random samples from the model. We can always condition on the representations learned by the model, but these will not be disentangled or interpretable.
Thus, OOD generation not only evaluates the ability of a model to compose learned concepts in new ways, but implicitly also the ability of a model to identify and capture underlying concepts of interest. Accordingly, we argue that OOD generation is a more appropriate evaluation metric.
While the numerical results on synthetic data above indicate better performance on OOD generation using our module, a more intuitive visual evaluation on benchmark datasets is provided in Section 5.2, Figure 5.2.
5.2 Large-scale models and benchmarks
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 4: Additional examples of concept composition in MNIST (left), 3DIdent (middle), and CelebA (right). In MNIST and 3DIdent, the samples are OOD: The training data did not contain any examples with these concepts composed together. The CelebA results are an ablation to understand the effect of conditioning vs intervention, and so there is some leakage between concepts, where the interventional datasets (MNIST, CelebA) show no leakage.
We now evaluate performance when our module is incorporated into a complex black-box model (NVAE of Vahdat and Kautz (2020)) on more challenging datasets, again including MNIST but now also 3DIdent (Zimmermann et al., 2021) and CelebA (Liu et al., 2015). Since CelebA does not have interventional data, we use this as an important ablation to test how well our module performs when used with conditional data (i.e. by conditioning on attributes in CelebA).
Table 3: Comparison of models (BPD = bits per dimension)
Dataset Type Black-box BPD
MNIST Observational NVAE 0.144
MNIST Interventional NVAE 0.149
3DIdent Observational NVAE 0.673
3DIdent Interventional NVAE 0.754
CelebA Observational NVAE 2.08
CelebA Conditional NVAE 2.13
Table 3 summarizes the results of these training runs, where “observational” means that the standard NVAE (i.e. without a context module) is trained on the entire datasets, whereas “interventional” means that the context module was used. We observe competitive reconstruction, with only slight degradation when the context module is attached to the NVAE decoder. More importantly, Figure 5.2 demonstrates generation capabilities enabled by our module. Each row corresponds to generated samples from the trained model in different contexts:
  1. 1.
    The first row shows generated observational samples;
  2. 2.
  3. 3.
For MNIST and 3DIdent, the final row of compositional generations is genuinely OOD in that the training data does not contain any examples where both concepts have been intervened upon. CelebA is trained using conditional as opposed to interventional samples as an ablation on the sensitivity to non-interventional data. Although there is a slight drop in OOD perceptual accuracy on CelebA, this is to be expected since the model was trained on conditional data as opposed to interventions. Finally, observe the last row of Figure 5.2, which demonstrates the OOD generation capability enabled by our module: these contexts did not appear in the training data but our module is able to sensibly generate them by composing and generalizing from the concepts it learned according to different data contexts.
Acknowledgements
The work by AM was supported by Novo Nordisk Foundation Grant NNF20OC0062897.

### v2

@@ 3.2 Details
We assume given a black-box encoder-decoder pair, denoted by $\operatorname{enc}(\mathbf{x})$ and $\operatorname{dec}(\mathbf{z})$, respectively. We modify the decoder by appending a context module to the decoder. The context module has three layers: A representation layer, an intervention layer, and an expressive layer.
@@
    The first representation layer, as the name suggests, learns to represent concepts by implicitly inverting their linear representations $\mathbf{c}_{j}=C_{j}\mathbf{e}$ from the embeddings of the encoder. This is a linear layer between $\mathbf{c}\to\mathbf{e}$ in (4).
@@
    The intervention layer embeds these concept representations into the reduced form SEM (3). Each concept has its own context where it has been intervened upon. A key part of this layer is how it can be used to learn and enforce interventional semantics through this shared SEM. This layer corresponds to the second layer $\boldsymbol{\varepsilon}\to\mathbf{c}$ in (4).
@@
Crucially, no part of this SEM is fixed or known—everything is trained end-to-end. In particular, we do not assume a known causal DAG or even a known causal order. This stands in contrast to previous work on causal generative models (e.g. Kocaoglu et al., 2017; Yang et al., 2021).
@@
Assume for now that the concepts are each one-dimensional with $\dim(\mathbf{c}_{j})=\dim(\boldsymbol{\varepsilon}_{j})=1$; extensions to multi-dimensional concepts are straightforward and explained below. The structural coefficients $\alpha_{kj}\in\mathbb{R}$ capture direct causal effects between concepts, with $\alpha_{kj}\neq 0$ indicating the presence of an edge $\mathbf{c}_{k}\to\mathbf{c}_{j}$. An intervention on the $j$th concept requires deleting each incoming edge; i.e., setting $\alpha_{\cdot j}=0$, updating the outgoing edges from $\boldsymbol{\varepsilon}_{j}$, as well as replacing $\boldsymbol{\varepsilon}_{j}$ with a new $\boldsymbol{\varepsilon}_{j}^{\prime}$, which results from updating the incoming edges to $\boldsymbol{\varepsilon}_{j}$ in the preceding layer. Thus, during training, when we intervene on $\mathbf{c}_{j}$, we zero out the row $\alpha_{\cdot j}$ while replacing the column $\alpha_{j\cdot}$ with a new column $\beta_{j\cdot}$ that is trained specifically for this interventional setting. The result is $d_{c}+1$ intervention-specific layers: $A_{0}$ for the observational setting, and $A_{1},\ldots,A_{d_{c}}$ for each interventional setting. This yields a three-way tensor where each slice of this tensor captures a different context that corresponds to intervening on different concepts. At inference time, to generate interventional samples we simply swap out the $A_{0}$ for $A_{j}$. Moreover, multi-target interventions can be sampled by zeroing out multiple rows and substituting multiple $\beta_{j\cdot}$’s and $\boldsymbol{\varepsilon}_{j}$’s.
@@
Everything above goes through if we allow additional flexibility with $\dim(\mathbf{c}_{j})>1$ and $\dim(\boldsymbol{\varepsilon}_{j})>1$, potentially even with $\dim(\mathbf{c}_{j})\neq\dim(\boldsymbol{\varepsilon}_{j})$. In practice, we implement this via two width parameters $0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ and $0pt_{c}=\dim(\mathbf{c}_{j})$. The design choice of enforcing uniform dimensionality is not necessary, but is made here since in our experiments there did not seem to be substantial advantages to choosing nonuniform widths. With these modifications, $\alpha_{kj}$ becomes a $0pt_{\varepsilon}\times 0pt_{c}$ matrix; instead of substituting rows and columns above, we now substitute slices in the obvious way. The three-way tensor $A$ is now $(d_{c}+1)\times 0pt_{\varepsilon}d_{c}\times 0pt_{c}d_{c}$-dimensional.
@@
An appealing aspect of this architecture is that it can be viewed as an approximation to an identifiable model over causally disentangled concepts. Formally, we assume given observations $\mathbf{x}$, concepts $\mathbf{c}$, and a collection of embeddings $\mathbf{e}$ such that $\mathbf{x}=f(\mathbf{e})$. Additionally assume a noisy version of the linear representation hypothesis, i.e., $\mathbf{c}_{j}=C_{j}\mathbf{e}+\varepsilon_{j}$ with $\varepsilon_{j}\sim\mathcal{N}(0,\Omega_{j})$. Then we have the following identifiability result:
Theorem 3.3 (Identifiability).
Assume that the rows of each $C_{j}$ come from a linearly independent set and $f$ is injective and differentiable. Then, given single-node interventions on each concept $\mathbf{c}_{j}$, the representations $C_{j}$ and latent concept distribution $p(\mathbf{c})$ are identifiable.
For a formal statement, see Theorem B.1 in Appendix B. As opposed to solving a causal representation learning problem, the causal semantics used in this paper have been ported into a generative model that give solutions to the problem of learning latent distribution structure, as in Theorem 3.3. We note also that the proof of Theorem 3.3 itself is nontrivial, and involves tracing the effects of interventions through the linear representations; since our focus is on implementation and practical aspects, we defer further discussion to Appendix B.
@@ 3.3 Summary of architecture
    Black-box. There is no coupling between the embedding dimension and the number of concepts, or the complexity of the SEM. This allows for arbitrary black-box encoder-decoder architectures to be used to learn embeddings, from which a causal model is then trained on top of. This can be trained end-to-end, fine-tuned after pre-training, or even partially frozen, allowing for substantial computational savings.
@@
    Causality. The intervention layer is a genuine causal model that provides rigorous causal semantics that allow sampling from arbitrary concept interventions, including interventions that have not been seen during training.
@@
    Identifiability. The architecture is based on an approximation to an identifiable concept model (Theorem 3.3), which provides formal justification for the intervention layer as well as reproducibility assurances.
@@
    Flexibility. The context module is itself arbitrarily flexible, so there is no risk of information loss in representing concepts with the embeddings $\mathbf{e}$. Of course, information loss is possible if we compress this layer too much (e.g., by choosing $d_{c}$, $k$, $0pt_{\varepsilon}$, or $0pt_{c}$ too small), but this is a design choice and not a constraint of the module itself.
Thus the only tradeoff between representational capacity and causal semantics is design-based: The architecture itself imposes no constraints. The causal model can be arbitrarily flexible and chosen independent of the encoder-decoder pair, which are also allowed to be arbitrary.
4 Empirical evaluation
[…]
@@ 4 Empirical evaluation
    Concept learning (Section 4.1): How well does the context module learn representations that correspond to each concept? To measure this, we generate interventional samples by intervening on a single concept, and directly compare these interventional samples to held-out samples using the sliced Wasserstein (SW) metric (18).
@@
    Composition (Section 4.2): How well does the context module learn to compose concepts together in novel ways that have not been seen during training? To measure this, we generate samples by intervening on multiple concepts, and directly compare these new samples to held-out samples using the SW metric. By design, interventions on multiple concepts are never seen during training, so this is genuinely OOD.
@@
    Reconstruction (Section 4.3): How well does the context module reconstruct samples from each context? This can be evaluated using standard measures of generation quality; we use bits per dimension (BPD) for comparison with prior work. Since this is computed across all contexts, this captures the model’s ability to reconstruct samples from diverse contexts.
@@
    Fine-tuning (Section 4.4): Can pre-trained models leverage the context module via fine-tuning to improve their learned representations and enable compositional generation? Instead of training end-to-end, we start with a pre-trained model and fine-tune it with the context module attached. This is evaluated on the concept learning, composition, and reconstruction tasks.
For concept learning and composition, including OOD generation (Section 2), we report the sliced Wasserstein (SW) distance (Bonneel et al., 2015; Flamary et al., 2021) between a held-out sample from the ground-truth distribution and a sample generated (not reconstructed) by the trained model. For reconstruction, we report the standard ELBO loss (Kingma and Welling, 2014) in bits per dimension.
Methods, ablations, baselines
[…]
Datasets
We compare these models to our context module on four datasets: 3DIdent (Zimmermann et al., 2021), Morpho MNIST (Castro et al., 2019), and two versions of a new semi-synthetic environment called quad—one with independent concepts and another with (causally) dependent concepts. While 3DIdent and Morpho MNIST are established baselines, quad is a novel visual environment with intervenable latents (colour, shape, size, orientation) that explicitly enables sampling from ground-truth OOD composed contexts. See Appendix C for more details on this dataset. We introduce quad in order to carefully evaluate model performance in a controlled environment over different random seeds (and hence assess statistical variability). Thus, our evaluations use quad to assess reproducibility and statistical significance, whereas 3DIdent and MNIST are more appropriate for testing our approach on downstream tasks such as image generation.
Due to space constraints, we report only a representative slice of the experiments here; full details and additional experiments can be found in Appendices CE. This includes detailed performance breakdowns across contexts; the effects of regularization and architectural capacity; and additional figures.
4.1 Concept learning
Table 1, in the Concept columns, shows a quantitative evaluation of concept learning on the quad, MNIST, and 3DIdent datasets. We draw three main conclusions from this table, by comparing our context module variants to each of the other methods (rows) in the ablations and baselines.
First, comparing the three CM rows to the ablations in Table 1 (specifically, base and base pooled), we see that the context module excels at learning individual concepts across all datasets and models, despite the base models achieving slightly better average reconstruction performance (see Section 4.3). Moreover, Figure 2 shows that there is essentially no perceptual difference between samples from the base models and the context module. Additionally, across multiple training runs over different random seeds, our context module consistently outperforms the base models across quad concepts—in this case the simpler model/data allows more replicates to be run to evaluate statistical significance and standard errors.
Second, comparing the CM rows to the base row of Table 1, we see that the context module outperforms the base model trained only on the observational (rather than pooled) data. This confirms that the context module is indeed successfully exploiting learned invariances across the different contexts, providing it with performance and computational resource advantages compared to models trained independently on the different contexts.
Table 1: Context module (CM) vs. fine-tuning, ablations, and baselines on three tasks across four datasets. Best performance per task/dataset in bold. Base model indicated by $^{*}=$ lightweight VAE, $^{**}=$ NVAE.
quad* (independent) quad* (dependent) MNIST** 3DIdent**
Concept Compo Recon Concept Compo Recon Concept Compo Recon Concept Compo Recon
ours CM (end-to-end) 0.063 0.081 0.519 0.116 0.127 0.828 0.035 - 0.135 0.014 0.051 0.528
CM (fine-tune) 0.059 0.086 0.518 0.112 0.121 0.840 0.037 - 0.136 0.031 0.063 0.549
CM (frozen) 0.086 0.132 0.559 0.121 0.134 0.889 0.052 - 0.161 0.048 0.063 0.609
ablations base 0.175 0.288 0.444 0.170 0.222 0.767 0.102 - 0.271 0.056 0.092 4.966
base pooled 0.166 0.248 0.446 0.167 0.212 0.746 0.105 - 0.139 0.049 0.056 0.559
CM pooled 0.202 0.306 0.840 0.166 0.203 0.752 0.099 - 0.143 0.041 0.068 0.516
baselines Ada-GVAE 0.335 0.355 2.017 0.375 0.377 2.443 0.500 - 1.035 0.249 0.251 3.267
BetaTCVAE 0.330 0.346 1.973 0.366 0.374 2.408 0.484 - 1.036 0.257 0.254 3.261
TVAE 0.334 0.341 2.048 0.370 0.368 2.455 0.486 - 1.037 0.254 0.256 3.282
Third, comparing the CM rows to the ablation in CM pooled—which pools all the data into a single context in order to circumvent the context module’s intervention mechanism while leaving the rest of the context module architecture intact—we see that CM consistently outperforms this ablation. This demonstrates that the increased performance of the context module is due our module’s intervention mechanism and its resulting ability to better leverage context-separated data, as opposed to spuriously benefiting from a different architecture or increased complexity.
Finally, for a qualitative evaluation, we look at Figure 2, in the Concept 1 and Concept 2 rows, to see concept learning in CM+NVAE across three different datasets. For example, looking at the 3DIdent results in Figure 2 (middle), the Observation row shows the learned observational distribution of images to contain background colours and object colours both in the orange–green range ( ). The Concept 1 row, produced by intervening on the learned concept of background colour, shows that the background colour is shifted to the blue–red range while other features, like object colour, remain invariant with the observational samples. Analogously, the Concept 2 row shows a shift of object colour to the blue–red range ( ) while the other features remain invariant. These illustrate that CM+NVAE successfully learned the concepts of background colour and object colour, in line with the notion of causal disentanglement in Definition 2.1.
4.2 Concept composition
Table 1, in the Compo columns, provides a quantitative evaluation of concept composition on the quad and 3DIdent datasets, while the Compo row of Figure 2 provides a qualitative evaluation across the MNIST and 3DIdent datasets. Since Morpho MNIST does not include ground truth compositional samples, we are not able to evaluate OOD generation on this dataset. In each evaluation, the model uses the intervention layer (Section 3.2) to intervene on pairs of learned concepts, i.e., to compose two concepts together. By holding out from training examples where concepts are composed, we are assured that the model has not seen any of these compositions. The only way to compose these concepts together is through the implicit causal model (via the reduced form SEM (3)) that is learned during training. We evaluate model performance with the SW metric (18), comparing concept compositions from the model against ground-truth hold out compositions.
The same patterns in performance hold here for concept composition as we saw for concept learning in Section 4.1—namely, our context module consistently achieves better performance than baselines and ablations, demonstrating successful concept composition due to the module’s intervention mechanism. For example, looking again at the 3DIdent results in Figure 2 (middle), we see that the colour range of both the background and the object are shifted from the orange–green range ( ) to the blue–red range ( ).
These experiments demonstrate the context module’s ability to learn causally disentangled representations through both concept learning and composition.
(a)      (b)
Figure 2: (a) Examples of concept learning and composition in MNIST (left) and 3DIdent (right) using CM (end-to-end), our context module augmenting the base NVAE—these images are generated, not reconstructed. The first row shows generated observational samples; The second and third rows show samples of learned concepts, generated by intervening on individual concepts, e.g., ‘Concept 1’ is ‘scaled’ for MNIST; The final row shows OOD composition, generated by simultaneously intervening on pairs of learned concepts. (b) Observational samples from the context module vs. the base model show no perceptual loss.
4.3 Reconstruction
The Recon columns of Table 1 compare the validation reconstruction performance of our approach against the alternatives. We draw two conclusions from these results. First, adding the context module incurs only a minor quantitative cost in reconstruction performance, while granting additional concept learning and composition abilities explored in Sections 4.1-4.2. Second, Figure 2 demonstrates that despite this small quantitative cost, the perceptual quality of the samples is still very high and indistinguishable from the base model. This validates the central motivation behind the context module, namely that it can imbue existing black-box models with additional concept learning and causal composition abilities without sacrificing perceptual metrics or downstream performance.
4.4 Fine-tuning
Finally, an important aspect of our method is that it need not be trained end-to-end but can instead be successfully fine-tuned from a frozen base mode. The second (CM (fine-tune)) and third (CM (frozen)) rows of Table 1 evaluate the context module on concept leaning and composition using two fine-tuning variants. Both variants start with a pre-trained model and then insert the context module and resume training, either only updating the context module weights in the frozen variant, or also fine-tuning the non-context module weights in the fine-tune variant.
[…]
5 Conclusion and Limitations
Designing generative models with structured latent spaces that enable intervention, composition, and OOD generation is an outstanding challenge. Existing approaches either sacrifice structure for flexibility, or flexibility for structure. We argue that such a tradeoff is not necessary and that arbitrarily flexible models can indeed learn useful structure.
To this end, we propose a simple context module that can be appended to a black-box decoder that learns causally disentangled latent spaces. The context module enables genuinely OOD, compositional generation without hardcoding prior knowledge or structure into the model, and it avoids the difficult problem of latent causal discovery.
Our experiments provide both quantitative and qualitative evidence for this. While existing evaluations have focused on reconstruction as a metric to quantify OOD performance, we argue that OOD generation is in fact more appropriate since it also measures the ability of a model to learn useful concepts that can be manipulated at inference time.
Since our module imposes no hard constraints, it may be useful in combination with other modules. Understanding how the context module interacts with other architectures is a natural next step and a promising direction for future work. We also encourage the community to explore alternative metrics for evaluating genuine OOD generalization.
Finally, understanding the cost and tradeoffs of variational approximations compared to, say, exact maximum likelihood estimation is important, especially in causal applications. Indeed, this can make estimating the exact causal graph much more challenging. This motivates our fundamental design principle: to use a “causal enough” architecture to obtain composability and OOD generation while otherwise leveraging already empirically proven non-causal architectures.
Acknowledgments